OpenStudy (anonymous):

FAN N MEDAL Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21

OpenStudy (anonymous):

j-14=m j+4=3(m+4) first equation reworked to j=14+m by adding 14 to both sides. plug that in for j in the second equation and you get 14+m+4=3(m+4). Simplify that and you get 18+m=3m+12. Simplified again is 2m=6 and ultimately m=3 Mary is 3, Jane is 17 (14 years older)

OpenStudy (anonymous):

You are correct :) thanks kind sir